dx/y^2 z=dy/xz=dz/yz solution ?


1
Expert's answer
2020-12-25T15:09:50-0500

"\displaystyle

\frac{\mathrm{d}x}{y^2 z} = \frac{\mathrm{d}y}{xz} = \frac{\mathrm{d}z}{yz}\\

\textsf{Comparing the first and the second equation,}\\
\frac{\mathrm{d}x}{y^2 z} = \frac{\mathrm{d}y}{xz}\\

\frac{\mathrm{d}x}{y^2} = \frac{\mathrm{d}y}{x}\\

x\mathrm{d}x - y^2\mathrm{d}y = 0\\

\int\,x\mathrm{d}x - \int\,y^2\mathrm{d}y = 0\\


\frac{x^2}{2} - \frac{y^3}{3} = C\\

\textsf{At}\,\, C = 0\\


\frac{x^2}{2} - \frac{y^3}{3} = 0,\,\, \frac{3x^2}{2} = y^3,\,\, y = \sqrt[3]{\frac{3x^2}{2}}\\

\textsf{Comparing the first and the third equation,}\\
\frac{\mathrm{d}x}{y^2 z} = \frac{\mathrm{d}z}{yz}\\

\frac{\mathrm{d}x}{y} = \mathrm{d}z\\

\frac{\mathrm{d}x}{\sqrt[3]{\frac{3x^2}{2}}} = \mathrm{d}z\\

\int\frac{\mathrm{d}x}{\sqrt[3]{\frac{3x^2}{2}}} = \int\mathrm{d}z\\

\int\,\, \sqrt[3]{\frac{2}{3}} x^{-\frac{2}{3}}\mathrm{d}x = \int\mathrm{d}z\\

\sqrt[3]{\frac{2}{3}} \cdot 3x^{\frac{1}{3}} + C = z\\


z = 2^{\frac{1}{3}} 3^{\frac{2}{3}} + C = 18^{\frac{1}{3}} \cdot x^{\frac{1}{3}} + C\\

z = \sqrt[3]{18x} + C\\


\therefore\phi\left(z - \sqrt[3]{18x},\,\, \frac{x^2}{2} - \frac{y^3}{3}\right) = 0\,\,\textsf{is a solution to the PDE}"